jigglypuff314 (jigglypuff314):

Suppose that the water level of the conical paper cup that is 15 cm and a 5 cm radius is rising at 4 cm/s. How fast is the water being poured into the cup when the water is 8 cm deep?

jigglypuff314 (jigglypuff314):

looking for dV/dt given dh/dt = 4 and h = 8 V = 1/3 pi r^2 h take derivative -> dV/dt = 2/3 pi r (dr/dt) (dh/dt) I'm stuck there because I don't know dr/dt :/